an aqueous solution contain 14.0 g of silver nitrate was added to an aqueous solution contain 4.83 g of calcium chloride. Assuming all of the silver nitrate is consumed, how many grams of calcium chloride would required to react with the silver nitrate? is there enough calcium chloride to consume all of the silver nitrate?
Answer #1

Molar mass of AgNO 3 = 107.87 + 14.0067 + ( 3 16.00) = 169.88 g / mol

Molar mass of CaCl 2 = 40.08 + ( 2 16.00 ) = 72.08 g / mol

Consider reaction, 2 AgNO 3 (aq) + CaCl 2 (aq) 2 AgCl (s) + Ca(NO 3 ) 2 (aq)

From reaction, 2 mol  AgNO 3 1 mol CaCl 2

i e 2 169.88 g  AgNO 3 72.08 g CaCl 2

14.0 g AgNO 3 72.08 14.0 / ( 2 169.88 )  g CaCl 2

14.0 g AgNO 3 2.97 g CaCl 2

For complete reaction of 14.0 g AgNO 3 , 2.97 g CaCl 2 is required.

ANSWER : 2.97 g CaCl 2